OpenStudy (anonymous):

between which two consecutive integers does square root 147 lie?

OpenStudy (anonymous):

find the perfect square under the number you want find (12x12=144) and the one just above which is (13x13=169) so 12 and 13 would be your answer

OpenStudy (anonymous):

thank you! :]

OpenStudy (mathmale):

I recognize that 12^2 = 144, so given 145, I'd assume that the square root of 147 lies in the interval \[(\sqrt{144},\sqrt{169}).\]Why choose 169? Because it's the next perfect square after 144. Why choose 144? Because it's the perfect square closest, but smaller than, 147.

OpenStudy (anonymous):

thank you for the detailed explanation!